OpenStudy (anonymous):

Can someone please help me with this derivative? Im really close but am missing something... d/dx (Tanh(3x))^1/2)

OpenStudy (anonymous):

\[\frac{ d }{dx}=(\tan 3x)^{\frac{ 1 }{ 2}}\] Is that it?

OpenStudy (anonymous):

tanh...hyperbolic function.

OpenStudy (anonymous):

OK, it's differentiated the same. So, using the chain rule, you'd do the outside first. \[\frac{ 1 }{ 2 }(\tanh(3x))^{\frac{- 1 }{ 2 }}\]

OpenStudy (anonymous):

Then the next layer in...\[\sec ^{2}(3x)\]

OpenStudy (anonymous):

And finally the last layer in... 3. When you put it all together you get \[\frac{ 3 sech ^{2}3x}{2\sqrt{\tanh 3x} }\]

OpenStudy (anonymous):

Got it thanks alot Becki!